OpenStudy (anonymous):

show that the expression u = sin (x-vt) describing a sinusoidal wave, satisfies the wave equation. Show that in general u = f(x - vt) and u = f(x + vt) satisfies the wave equations, where f is any function with a second derivative

OpenStudy (anonymous):

@mukushla

OpenStudy (anonymous):

slow down buddy . you wrote u=f(x-vt) twice

OpenStudy (anonymous):

i wrote once, u = f(x - vt) and u = f(x + vt)

OpenStudy (anonymous):

The wave equation is: \[\frac{\partial^2 u}{\partial t^2} = v^2\frac{\partial^2 u}{\partial x^2}\] In your case, \(u = \sin(x - vt)\) And \[u_{tt} = -\sin(x - vt) * v^2 \] \[u_{xx} = -\sin(x - vt)\] So since \(u_{tt} = v^2 u_{xx}\), u satifies the wave equation

OpenStudy (anonymous):

good job @Meepi i appreciate

OpenStudy (anonymous):

wait a sec

OpenStudy (anonymous):

isn't it 1/v2 ??

OpenStudy (anonymous):

No, \(u_{tt} = -\sin(x - vt) * v^2\) and \(u_{xx} = -\sin(x - vt)\) the v^2 is from the chain rule when you derive with respect to t in \(u_{tt}\) since \(u_{xx} = -\sin(x - vt)\), \(u_{tt} = u_{xx} * v^2\)

OpenStudy (anonymous):

isn't it \[\frac{ ∂^{2}u }{ ∂t^{2} } = \frac{ 1 }{ v^{2} } \frac{ ∂^{2}u }{ ∂x^{2} }\] ??? you typed \(\frac{ ∂^{2}u }{ ∂t^{2} } = v^{2} \frac{ ∂^{2}u }{ ∂x^{2} } \)

OpenStudy (anonymous):

i thought \(\frac{ ∂^{2}u }{ ∂t^{2} } = \frac{ 1 }{ v^{2} } \frac{ ∂^{2}u }{ ∂x^{2} }\)

OpenStudy (anonymous):

ohh my bad.., sorry

OpenStudy (anonymous):

\[\frac{ ∂^{2}u }{ ∂x^{2} } = \frac{ 1 }{ v^{2} } \frac{ ∂^{2}u }{ ∂t^{2} }\]

OpenStudy (anonymous):

It's alright your text book probably has the partial derivatives swapped

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ur allright :), cool thanks :)

OpenStudy (anonymous):

Also, for the general case, just use the chain rule as well to get \[u_{xx} = f''(x + vt)\] \[u_{tt} = f''(x + vt) * v^2\] Then \[u_{xx} = \frac{u_{tt}}{v^2}\]

OpenStudy (anonymous):

Same can be done for u = f(x - vt)

OpenStudy (anonymous):

ok.., i got it now :)

OpenStudy (anonymous):

awesome :D

OpenStudy (anonymous):

wanna help me with Legendre series ??

OpenStudy (anonymous):

Haven't really done much of that but I can give it a shot :)

OpenStudy (anonymous):

go to http://openstudy.com/study#/updates/5172d293e4b0f872395bc5be